Twice the difference of a number and 6 is equal to three times the sum of the number and 4. Find the number.
Answer by rapaljer(4551)

Let x = the number

2(x-6)=3(x+4)
2x-12=3x+12

Subtract 2x from each side to get all the variables on the right side:
2x-2x -12 = 3x-2x +12
-12 =X+12

Subtract 12 from each side to get all the number terms on the OPPOSITE side.
-12-12=x+12-12
-24 = x

R^2